OpenStudy (anonymous):

solve 3(cscx-2)=2-cscx

OpenStudy (anonymous):

3 cos(2x) − cos x + 2 = 0 3 (2 cos²x − 1) − cos x + 2 = 0 6 cos²x − 3 − cos x + 2 = 0 6 cos²x − cos x − 1 = 0 (3 cosx + 1) (2 cosx − 1) = 0 3 cosx + 1 = 0 cosx = −1/3 x = arccos(−1/3), 360 − arccos(−1/3) x = 109.47°, 250.53° 2 cosx − 1 = 0 cosx = 1/2 x = 60°, 300

OpenStudy (anonymous):

hope's it helps

OpenStudy (anonymous):

thannk you!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

plz medal me plz